# Prompting Large Language Models for Recommender Systems: A Comprehensive Framework and Empirical Analysis

### v1

@@ 3.3.2. Configuration and Implementation
As for ProLLM4Rec, we can directly evaluate the cold-start recommendation ability of large language models in the zero-shot setting, as well as evaluate the fine-tuned ability with few or full recommendation samples in the fine-tuning setting. Considering the typical scenarios of LLMs in recommender systems, we conduct experiments on two representative task settings, i.e., (1) the zero-shot ranking task without modifying parameters of LLMs (Dai et al., 2023a; Hou et al., 2023; Gao et al., 2023; Ma et al., 2023) and (2) the click-through rate prediction task with LLMs tuned (Bao et al., 2023b; Kang et al., 2023; Fu et al., 2023; Shi et al., 2023).
(a) Zero-shot ranking task (Hou et al., 2023)
(b) Fine-tuned CTR prediction task (Bao et al., 2023b)
Figure 2. The basic prompts used in the MovieLens-1M dataset for the following experiments.
@@
$\bullet$ Zero-shot ranking task. On the one hand, we evaluate the zero-shot recommendation performance of LLMs for cold-start scenarios to study the effect of LLMs and the design of prompts. In this paper, our approach mainly concentrates on ranking tasks that better reflect the capabilities of LLMs (Dai et al., 2023a; Hou et al., 2023; Gao et al., 2023; Ma et al., 2023). As shown in Fig. 1, information of users and items is encoded into the prompt as inputs for LLMs. In this setting, we do not modify the parameters of LLMs, so the evaluated models are closed-source large language models or open-source models without fine-tuning. To conduct experiments on the impact of each factor on recommendation results with LLMs, we implement the overall architecture based on the open-source recommendation library RecBole (Zhao et al., 2021, 2022; Xu et al., 2023) and the zero-shot re-ranker LLMRank (Hou et al., 2023). Our basic prompt used in the MovieLens-1M dataset for the zero-shot ranking task is shown in Fig. 2(a).
@@
$\bullet$ Click-through rate prediction task with LLMs tuned. On the other hand, we evaluate the fine-tuned recommendation performance of LLMs to explore how LLMs adapt to recommendation scenarios with data provided (Bao et al., 2023b; Kang et al., 2023; Fu et al., 2023; Shi et al., 2023). Although our framework can be generalized to various recommendation tasks, we concentrate on exploring the fine-tuning performance of LLMs with point-wise Click-Through Rate (CTR) prediction tasks to reduce selection bias. In this setting, we not only consider fine-tuning LLMs using recommendation data, but also devise a two-stage approach of using instruction data to fine-tune LLMs first, and then implement recommendation fine-tuning for further adaptation. Specifically, we compare the fine-tuned recommendation performance of the original LLM and the LLM after instruction tuning, respectively. LLaMA-7B (Touvron et al., 2023a) and LLaMA2-7B (Touvron et al., 2023b) are the original models, while Alpaca-lora-7B (Taori et al., 2023) and LLaMA2-chat-7B (Touvron et al., 2023b) are LLMs after instruction tuning. In terms of the tuning strategies of LLMs, we report results with both parameter-efficient fine-tuning and complete fine-tuning. We implement the fine-tuning framework based on the open-source library transformers and the instruction tuning code of LLaMA with Stanford Alpaca data  [^2] The repository of Alpaca-LoRA: https://github.com/tloen/alpaca-lora.. Our instruction data used in the MovieLens-1M dataset for the click-through rate prediction task is illustrated in Fig. 2(b).
@@ 3.3.3. Evaluation Metrics
As for the zero-shot ranking task, considering economic and efficiency factors, we refer to existing literature (Hou et al., 2023; Shu et al., 2023) to randomly sample 200 users instead of evaluating results on the whole dataset. For each user from the sample set, we sort all items that the user interacts with in the chronological order. Then, we evaluate the results based on the leave-one-out strategy and treat the last interacted item as the ground truth. For performance comparison, we fix the length of candidate items to 20 as in (Hou et al., 2023), and mix the other 19 items from the ground truth in random positions by default. As for evaluation, we utilize two widely used ranking metrics in recommender systems, i.e., Recall (Buckland and Gey, 1994) and Normalized Discounted Cumulated Gain (NDCG) (Järvelin and Kekäläinen, 2002). Since 20 items are selected for candidate generation, we set $k=20$ for Recall@$k$ to measure whether all ground-truth items have been recalled. In existing literature, the re-generation method of multiple results until the format requirements are met can be employed to obtain the final response (Li et al., 2023i), while we consider the recall capability within one inference for fair evaluation. Furthermore, we set $k=1,10,20$ for NDCG to explore the detailed recommendation performance in terms of ranking abilities. To assure a scientific research, we repeat each experiment three times and take the average value as the final result.
As for the CTR prediction task, we first sort the original dataset by timestamp, use the latest 10,000 records for training and evaluation, and regard other previous data as the interaction history of users. Then, we split the ten thousand interactions into the training, validation and test sets in ratio of 8:1:1. For each interaction, we retain 10 historical interacted items as user representations, and set thresholds based on rating data to obtain user preferences (Bao et al., 2023b). Interactions with a rating higher than or equal to the threshold will be considered items that the user likes, while the opposite indicates dislikes. For the MovieLens-1M dataset, the rating threshold is 4, and we set 5 for the Amazon-Books dataset. We employ the training data to fine-tuning LLMs and evaluate the recommendation performance on the test set. The validation set is used for selecting best checkpoints during the training process. As for evaluation of prediction, we utilize the widely used metric for CTR predictions, i.e., accuracy. In a random state, the accuracy is around 0.5.
@@ 3.3.4. Discussion on Variable Factors
To conduct our analysis, we mainly focus on two key aspects, i.e., LLMs and prompts. As for the effects of LLMs, we analyze the impact of public availability, tuning strategies, model architecture, parameter scale, and context length on recommendation results based on the classification of LLMs. As for prompt engineering, we further analyze the impact of four important components of prompts, i.e., task description, user interest modeling, candidate items construction and prompting strategies. Due to the fact that all factors have an impact on the final result, it is also crucial to select the other aspects when focusing on one aspect. Limited by resources and efficiency, it is neither necessary nor feasible to exhaust all possibilities. When not explicitly specified, the LLM uses ChatGPT released in June 2023 to ensure the consistent recommendation quality. For the design of prompts, we refer to the template in (Hou et al., 2023) and emphasize the most recent items to re-rank the random 20 candidate items. The default method for modeling user interest is to concatenate the title sequence of recently interacted items using natural languages.
@@ 4. THE IMPACT OF LARGE LANGUAGE MODELS AS RECOMMENDER SYSTEMS
LLMs are the core of our framework ProLLM4Rec, and directly determine the performance of LLMs as recommender systems (Liu et al., 2023d, c; Kang et al., 2023; Hou et al., 2023). Therefore, it is worth exploring how to choose a suitable LLM as the foundation for recommendation. In this section, we compare the differences between LLMs and traditional models on the recommendation performance, discuss how the different properties and tuning strategies of LLMs affect the recommendation results, present the limitations of LLMs as recommenders, and draw empirical conclusions through systematic experiments.
4.1. Classification of Large Language Models
In this paper, we consider language models that have a size larger that one billion as LLMs. In line with existing researches (Zhao et al., 2023; Wu et al., 2023b; Lin et al., 2023a; Li et al., 2023l; Fan et al., 2023), LLMs can be categorized into different classes from several perspectives. Most typically, LLMs can be divided into open-source and closed-source models in terms of the public availability. When it comes to leveraging LLMs as the foundation model in recommender systems, tuning strategies can adjust LLMs towards specific recommendation tasks. From the perspective of the model architecture, various LLMs can also be categorized into types of encoder-decoder, causal decoder, and prefix decoder (Zhao et al., 2023). For the same framework of a LLM, it is widely acknowledged that the parameter scale and context length are two key factors that jointly determine abilities of LLMs (Lin et al., 2023b; Hou et al., 2023). To explore the recommendation performance with respect to different variants of LLMs, we focus on five aspects, i.e., public availability, tuning strategies, model architecture, parameter scale and context length as follows.
@@ 4.1.1. Public Availability
According to whether the model checkpoints can be publicly obtained, existing LLMs can be divided into open-source models and closed-source models, and both can be leveraged as recommender systems.
$\bullet$ Open-source models refer to LLMs whose model checkpoints can be publicly accessible. As shown in Table 2, researchers often use recommendation data to fine-tune open-source models for performance improvement. As the representative of open-source models, LLaMA (Touvron et al., 2023a) and its variants like Vicuna (Chiang et al., 2023) are widely used when leveraging LLMs for recommender systems (Zheng et al., 2023; Li et al., 2023b). Parameter-Efficient Fine-Tuning (PEFT) strategies such as Low-Rank Adaptation (LoRA) (Hu et al., 2021) are frequently adopted for recommendation data considering the trade-off between effect and efficiency (Liao et al., 2023; Bao et al., 2023b). Other models such as the Flan-T5 (Chung et al., 2022) series from Google Inc. and ChatGLM (Zeng et al., 2022) from Tsinghua University are also popular in the field of recommender systems. The publicly available checkpoints of open-source models provide flexibility for LLMs to modify parameters tailored for recommendation tasks.
$\bullet$ Closed-source models refer to LLMs whose model checkpoints can not be publicly accessible. For the closed-source LLMs utilized for recommenders, researchers generally study the zero-shot recommendation ability in cold-start scenarios. The most typical closed-source model is the ChatGPT series from OpenAI. The subsequent GPT-4 has stronger capabilities compared to ChatGPT, but it is still not open-source. In this paper, ChatGPT refers to the API function of “gpt3.5-turbo-4k-0613” unless specified. Without checkpoints, OpenAI provides several ways for researchers and users to improve the model performance on specific tasks, such as plugins for website browsing (Lewis et al., 2020) and interfaces for fine-tuning ChatGPT (Li et al., 2023h). However, the flexibility of closed-source models as recommender systems is still limited due to the expensive price and black-box parameters. Faced with this challenge, existing literature has explored to inject knowledge of recommender systems into closed-source models by means of prompt design (Yao et al., 2023; Wang and Lim, 2023), retrieval enhancement (Huang et al., 2023; Wang et al., 2023c; Shu et al., 2023) and combination of traditional recommendation models (Wang et al., 2023h; Gao et al., 2023).
@@ 4.1.2. Tuning Strategies
During the deployment of LLMs in recommender systems, we can also classify existing work depending on whether LLMs are fine-tuned, as well as the various fine-tuning strategies employed, i.e., not-tuning setting, fine-tuning setting and instruction tuning setting.
@@
$\bullet$ Not-tuning setting means evaluating the zero-shot recommendation ability of LLMs, which is generally used for closed-source models such as ChatGPT. By designing different prompt templates, LLMs without fine-tuning can be directly used for recommendation tasks such as click-through rate predictions (Di Palma et al., 2023; Kang et al., 2023), sequential recommendation (Hou et al., 2023; Wang et al., 2023c), and conversational recommender systems (He et al., 2023; Wang et al., 2023h; Jin et al., 2023). In this case, the user interest is expressed explicitly (e.g., ratings and reviews) or implicitly (interacted items of users), and the limited candidate items can be recalled by traditional models, while prompting strategies such as role prompting and chain of thoughts are used. Inspired by Artificial Intelligence Generated Content (AIGC), it is worth noting that the excellent generation ability of LLMs provides opportunities for generative recommendation (Liu et al., 2023a; Wang et al., 2023e). Without providing candidate items, generative language models can directly generate the desired items that users need based on recommendation requirements, and they can also be generalized into our framework ProLLM4Rec as shown in Table 2.
@@
$\bullet$ Fine-tuning setting means that using recommendation data to fine-tune LLMs as recommender systems. Considering cost and efficiency, researchers often use parameter-efficient fine-tuning (e.g., Low-Rank Adaptation of Large Language Models, LoRA (Hu et al., 2021)) to quickly adapt to recommendation scenarios (Bao et al., 2023b; Liao et al., 2023; Zheng et al., 2023). As for LLMs, open-source models based on LLaMA (Touvron et al., 2023a) are widely used, including but not limited to LLaMA, LLaMA2 and Vicuna (Chiang et al., 2023) with different parameter sizes. Based on whether candidate items are provided, existing work on fine-tuning LLMs can be further divided into two kinds of recommendation tasks. On the one hand, researchers have explored fine-tuning models for recommendation that provide candidate items such as rating, re-ranking and predictions. Specifically, the fine-tuning interface of the closed-source model ChatGPT has brought new breakthroughs to the research of LLMs, and there have been attempts to fine-tune ChatGPT for recommendation tasks (Li et al., 2023h). On the other hand, LLMs can also be fine-tuned for recall tasks of recommender systems by retrieving candidates from the whole item pool (Bao et al., 2023a; Lin et al., 2023c; Petrov and Macdonald, 2023; Li et al., 2023b). Through well-designed indexing, alignment, and retrieval strategies, directly generating recommendation items without providing lengthy candidate sequences is more suitable for practical application scenarios, which has not yet been fully explored.
@@
$\bullet$ Instruction tuning setting means providing template instructions of recommenders as prompts to tuning targeted LLMs, generally involving multiple recommendation tasks (Zhang et al., 2023f; Geng et al., 2022; Qiu et al., 2023; Chu et al., 2023). However, some fine-tuning methods (e.g., TALLRec (Bao et al., 2023b)) also involve the form of instructions. In order to avoid ambiguity, we consider the instructions of a single task template as fine-tuning settings, and the instructions of multiple tasks as instruction-tuning settings. As the backbone of recommender systems, researchers generally use T5 or Flan-T5 for various recommendation scenarios such as rating, ranking, retrieving, explanation generation and news recommendation. Besides Flan-T5, RecRanker (Luo et al., 2023b) is also proposed to instruction tuning LLaMA2 as the ranker for top-$k$ recommendation. By instantiating appropriate instructions for each recommendation task, this setting can also be summarized into our framework.

### v2

@@ 3.3.2. Configuration and Implementation
As for LLM-RS, we can evaluate the cold-start recommendation ability of LLMs in the zero-shot setting, as well as evaluate the fine-tuned performance with a few or full recommendation samples in the fine-tuning setting. Considering the typical scenarios of LLMs as recommender systems, we conduct experiments on two representative task settings, i.e., (1) the zero-shot ranking task without modifying parameters of LLMs (generative recommendation) (Dai et al., 2023a; Hou et al., 2024; Gao et al., 2023; Ma et al., 2023); and (2) the Click-Through Rate (CTR) prediction task with LLMs tuned (discriminative recommendation) (Bao et al., 2023b; Kang et al., 2023; Fu et al., 2023; Shi et al., 2024).
(a) Zero-shot ranking task (Hou et al., 2024).
(b) Fine-tuned CTR prediction task (Bao et al., 2023b).
Figure 2. Basic prompts for experiments on the MovieLens-1M dataset in LLM-RS. We explore two application scenarios: (1) prompting LLMs without tuning, and (2) fine-tuning LLMs as recommender systems. The experiments focus on sequential re-ranking and Click-Through Rate (CTR) prediction tasks, respectively.
@@
$\bullet$ Zero-shot ranking task. On the one hand, we evaluate the zero-shot recommendation performance of LLMs for cold-start scenarios to study the effect of LLMs and the design of prompts. In this paper, our approach mainly concentrates on ranking tasks that better reflect the capabilities of LLMs (Dai et al., 2023a; Hou et al., 2024; Gao et al., 2023; Ma et al., 2023). As shown in Fig. 1, information of users and items is encoded into the prompt as inputs for LLMs. In this setting, we do not modify the parameters of LLMs, so the evaluated models are closed-source LLMs or open-source models without fine-tuning. To conduct experiments on the impact of each factor on recommendation results with LLMs, we implement the overall architecture based on the open-source recommendation library RecBole (Zhao et al., 2021, 2022; Xu et al., 2023) and the zero-shot re-ranker LLMRank (Hou et al., 2024). Our basic prompt used in the MovieLens-1M dataset for the zero-shot ranking task is shown in Fig. 2(a).
@@
$\bullet$ CTR prediction task with LLMs tuned. On the other hand, we evaluate the fine-tuned recommendation performance of LLMs to explore how LLMs adapt to recommendation scenarios with data provided (Bao et al., 2023b; Kang et al., 2023; Fu et al., 2023; Shi et al., 2024). Although our framework can be generalized to various recommendation tasks, we concentrate on exploring the fine-tuning performance of LLMs with point-wise CTR prediction tasks to reduce selection bias. In this setting, we not only consider fine-tuning LLMs using recommendation data from scratch, but also devise a two-stage approach of using instruction data to fine-tune LLMs first, and then implement recommendation fine-tuning for further adaptation. Specifically, we compare the fine-tuned recommendation performance of the original LLM and the LLM after instruction tuning, respectively. As for the selection of LLMs, LLaMA-7B (Touvron et al., 2023a), LLaMA2-7B (Touvron et al., 2023b) and LLaMA3-8B (Dubey et al., 2024) are the original models, while Alpaca-lora-7B (Taori et al., 2023) and LLaMA2-chat-7B (Touvron et al., 2023b) are LLMs after instruction tuning. In terms of the tuning strategies of LLMs, we report results with both Parameter-Efficient Fine-Tuning (PEFT) and full-model fine-tuning. We implement the fine-tuning framework based on the open-source library transformers and the instruction tuning code of LLaMA with Stanford Alpaca data  [^2] The repository of Alpaca-LoRA: https://github.com/tloen/alpaca-lora.. Our basic prompts to fine-tune LLMs for the CTR prediction task on the MovieLens-1M dataset is illustrated in Fig. 2(b).
@@ 3.3.3. Evaluation Metrics
As for the zero-shot ranking task, considering economic and efficiency factors, we refer to existing literature (Hou et al., 2024; Shu et al., 2023) to randomly sample 200 users instead of evaluating results on the whole dataset. The average result of 200 users is the final indicator for performance comparison. For each user from the sample set, we sort all items that the user interacts with in the chronological order. Then, we evaluate the results based on the leave-one-out strategy and treat the last interacted item as the ground truth. For performance comparison, we fix the length of candidate items to 20 as in (Hou et al., 2024), and mix the other 19 items from the ground truth in random positions by default. As for evaluation, we utilize two widely used ranking metrics in recommender systems, i.e., Recall (Buckland and Gey, 1994) and Normalized Discounted Cumulated Gain (NDCG) (Järvelin and Kekäläinen, 2002). Since 20 items are selected for candidate generation, we set $k=20$ for Recall@$k$ to measure whether all ground-truth items have been recalled, and the metric Recall@$20$ can also be expressed as Coverage. In existing literature, the re-generation method of multiple results until the format requirements are met can be employed to obtain the final response (Li et al., 2023h), while we consider the recall capability within one inference for fair evaluation. Furthermore, we set $k=1,10,20$ for NDCG to explore the detailed recommendation performance in terms of ranking abilities. For a scientific research, we repeat each experiment three times and take the average value as the results.
As for the CTR prediction task, we first sort the original dataset by timestamp, use the latest 10,000 records for training and evaluation, and regard other previous data as the interaction history of users. Then, we split the interactions into the training, validation and test sets in ratio of 8:1:1. For each interaction, we retain 10 historical interacted items as user representations, and set thresholds based on rating data to obtain user preferences (Bao et al., 2023b). Interactions with a rating higher than or equal to the threshold will be considered items that the user likes, while the opposite indicates dislikes. For the MovieLens-1M dataset, the rating threshold is 4, and we set 5 for the Amazon Books dataset. We employ the training data to fine-tune LLMs and evaluate the recommendation performance on the test set. The validation set is used for selecting best checkpoints during the training process. As for evaluation of prediction, we utilize the widely used metric for CTR predictions, i.e., accuracy. In a random state, the accuracy is around 0.5.
@@ 3.3.4. Discussion on Variable Factors
To conduct our analysis for LLM-RS, we mainly focus on using LLMs as recommender systems with two key aspects, i.e., LLMs and prompts. As for the effects of LLMs, we analyze the impact of public availability, tuning strategies, model architecture, parameter scale, and context length on recommendation results based on the classification of LLMs. As for prompt engineering, we further analyze the impact of four important components of prompts, i.e., task description, user interest modeling, candidate items construction and prompting strategies. Due to the fact that all factors have an impact on the final result, it is also crucial to select the other aspects when focusing on one aspect. Limited by resources and efficiency, it is neither necessary nor feasible to exhaust all possibilities. When not explicitly specified, the LLM uses ChatGPT released in June 2023 to ensure the consistent recommendation quality. For the design of prompts, we refer to the template in (Hou et al., 2024) and emphasize the most recently interacted items to re-rank 20 candidate items. The default method for modeling user interest is to concatenate the title sequence of recently interacted items using natural languages.
@@ 4. THE IMPACT OF LARGE LANGUAGE MODELS AS RECOMMENDER SYSTEMS
LLMs are the core of our framework LLM-RS, and largely determine the performance of LLMs as recommender systems (Liu et al., 2023d, c; Kang et al., 2023; Hou et al., 2024). Therefore, it is worth exploring how to choose a suitable LLM as the foundation for recommendation. In this section, we compare the differences between LLMs and traditional models on the recommendation performance, discuss how the different properties and tuning strategies of LLMs affect the recommendation results, present the limitations of LLMs as recommenders, and draw empirical conclusions through systematic experiments.
4.1. Classification of LLMs
In this paper, we consider language models that have a size larger that one billion as LLMs. In line with existing researches (Zhao et al., 2023; Wu et al., 2024b; Lin et al., 2023a; Li et al., 2024d; Zhao et al., 2024), LLMs can be categorized into different classes from several perspectives. Most typically, LLMs can be divided into open-source and closed-source models in terms of the public availability. When it comes to leveraging LLMs as the foundation model in recommender systems, tuning strategies can adjust LLMs towards specific recommendation tasks. From the perspective of the model architecture, various LLMs can also be categorized into types of encoder-decoder, causal decoder, and prefix decoder (Zhao et al., 2023). For the same framework of a LLM, it is widely acknowledged that the parameter scale and context length are two key factors that jointly determine capabilities of LLMs (Lin et al., 2024a; Hou et al., 2024). To explore the recommendation performance w.r.t. different variants of LLMs, we focus on five aspects, i.e., public availability, tuning strategies, model architecture, parameter scale and context length as follows.
@@ 4.1.1. Public Availability
According to whether the model checkpoints can be publicly obtained, existing LLMs can be divided into open-source models and closed-source models. In what follows, we will explain that both the two categories of LLMs can be leveraged as recommender systems.
$\bullet$ Open-source models refer to LLMs whose model checkpoints can be publicly accessible. As shown in Table 2, researchers often use recommendation data to fine-tune open-source models for performance improvement. As the representative of open-source models, LLaMA (Touvron et al., 2023a) and its variants like Vicuna (Chiang et al., 2023) are widely used when leveraging LLMs for recommender systems (Zheng et al., 2024; Li et al., 2023c). Parameter-Efficient Fine-Tuning (PEFT) strategies such as Low-Rank Adaptation (LoRA) (Hu et al., 2021) are frequently adopted for recommendation data considering the trade-off between effect and efficiency (Liao et al., 2023; Bao et al., 2023b). Other models such as the Flan-T5 (Chung et al., 2024) series from Google Inc., ChatGLM (AI, 2024) from Zhipu AI and Qwen (Team, 2024) from Alibaba are also popular in the field of recommender systems. The publicly available checkpoints of open-source models provide flexibility for LLMs to modify parameters tailored for recommendation tasks.
$\bullet$ Closed-source models refer to LLMs whose model checkpoints can not be publicly accessible. For the closed-source LLMs leveraged as recommender systems, researchers generally study the zero-shot recommendation ability in cold-start scenarios. The most typical closed-source model is the ChatGPT series from OpenAI. The subsequent GPT-4 (OpenAI, 2023) and GPT-4o (OepnAI, 2024) have stronger capabilities compared to ChatGPT, but they are still not open-source. In this paper, ChatGPT refers to the API function of “gpt3.5-turbo-4k-0613” unless specified. Without checkpoints, OpenAI provides several ways for researchers and users to improve the model performance on specific tasks, such as plugins for website browsing (Lewis et al., 2020) and interfaces for fine-tuning ChatGPT (Li et al., 2023h). However, the flexibility of closed-source models as recommender systems is still limited due to the expensive price and black-box parameters. Faced with this challenge, existing literature has explored to inject knowledge of recommender systems into closed-source models by means of prompt design (Yao et al., 2023; Wang and Lim, 2023), retrieval enhancement (Huang et al., 2023; Wang et al., 2024c; Shu et al., 2023) and combination of traditional recommendation models (Wang et al., 2023f; Gao et al., 2023).
@@ 4.1.2. Tuning Strategies
During the deployment of LLMs in recommender systems, we can also classify existing work depending on whether LLMs are fine-tuned, as well as the various fine-tuning strategies employed, i.e., prompting without tuning, fine-tuning and instruction tuning.
@@
$\bullet$ Prompting without tuning means evaluating the zero-shot recommendation ability of LLMs, which is generally used for closed-source models such as ChatGPT. By designing prompt templates, LLMs without fine-tuning can be used for recommendation tasks such as CTR predictions (Di Palma et al., 2023; Kang et al., 2023), sequential recommendation (Hou et al., 2024; Wang et al., 2024c), and Conversational Recommender Systems (CRS) (He et al., 2023; Wang et al., 2023f; Jin et al., 2023). In this case, the user interest is expressed explicitly (e.g., ratings and reviews) or implicitly (interacted items of users), and the limited candidate items can be recalled by traditional models, while prompting strategies such as In-Context Learning (ICL) and Chain-of-Thoughts (CoT) are used. Inspired by Artificial Intelligence Generated Content (AIGC), it is worth noting that the excellent generation ability of LLMs provides opportunities for generative recommendation (Liu et al., 2023a; Wang et al., 2023c). Without providing candidate items, generative language models can generate the desired items that users need based on recommendation requirements, which can also be generalized into our framework.
@@
$\bullet$ Fine-tuning LLMs means that using recommendation data to fine-tune LLMs as recommender systems. Considering cost and efficiency, researchers often use parameter-efficient fine-tuning (e.g., LoRA (Hu et al., 2021)) to adapt to recommendation scenarios (Bao et al., 2023b; Liao et al., 2023; Zheng et al., 2024). As for LLMs, open-source models based on LLaMA (Touvron et al., 2023a) are widely used, including but not limited to LLaMA, LLaMA2 (Touvron et al., 2023b), LLaMA3 (Dubey et al., 2024) and Vicuna (Chiang et al., 2023) with different parameter sizes. Based on whether candidate items are provided, existing work on fine-tuning LLMs can be further divided into two kinds of recommendation tasks. On the one hand, researchers have explored fine-tuning models for recommendation that provide candidate items such as rating, re-ranking and CTR predictions. Specifically, the fine-tuning interface of the closed-source model ChatGPT has brought new breakthroughs to the research of LLMs, and there have been attempts to fine-tune ChatGPT for recommendation tasks (Li et al., 2023h). On the other hand, LLMs can also be fine-tuned for recall tasks of recommender systems by retrieving candidates from the whole item pool (Bao et al., 2023a; Lin et al., 2023b; Petrov and Macdonald, 2023; Li et al., 2023c). Through well-designed indexing, alignment, and retrieval strategies, generating recommendation items without providing lengthy candidate sequences is suitable for practical application scenarios (Hua et al., 2023; Chen et al., 2024a).
@@
$\bullet$ Instruction tuning means providing template instructions of recommenders as prompts to tuning targeted LLMs, generally involving multiple recommendation tasks (Zhang et al., 2023e; Geng et al., 2022; Qiu et al., 2023; Chu et al., 2023; Tan et al., 2024). However, some fine-tuning methods (e.g., TALLRec (Bao et al., 2023b)) also involve the form of instructions. In order to avoid ambiguity, we consider the instructions of a single task template as fine-tuning settings, and the instructions of multiple tasks as instruction-tuning settings. As the backbone of recommender systems, researchers generally use T5 (Raffel et al., 2020) or Flan-T5 (Chung et al., 2024) for various recommendation scenarios such as rating, ranking, retrieving, explanation generation and news recommendation. Besides Flan-T5, RecRanker (Luo et al., 2023) is also proposed to instruction tuning LLaMA2 (Touvron et al., 2023b) as the ranker for top-$k$ recommendation. By instantiating appropriate instructions for each recommendation task, this setting can also be summarized into our framework.